# Lanczos-Pascal approach to correlation functions in chaotic quantum systems

### v2

@@ Abstract
Introduction. Understanding out-of-equilibrium quantum dynamics is one of the most fundamental and challenging questions in the field of statistical physics. It has received significant attentions, from both theoretical and experimental perspectives [1, 2, 3, 4]. Despite the substantial progress that has been made, many fundamental aspects of the problem remain unresolved, and it continues to be the focus of intensive research in the field.
@@
To characterize the evolution of correlation functions, on the analytical side, except for trivial cases such as free fermions, general analytical predictions remain a major difficulty. In systems with conserved charge, hydrodynamics has been proven to be a useful tool to predict the long-time behavior of autocorrelation function [9, 10, 11]. However, it remains an assumption, and the extent of its applicability has yet to be fully established.
@@
Another approach, the recursion method [22, 23, 24], significantly regained attention thanks to the recent Operator Growth Hypothesis (OGH) [25]. The latter constitutes a statement on the so-called Lanczos coefficients of an observable; quantities that fully characterize its dynamics in the thermodynamic limit. Eventually code bases such as [26] position this method along other well-established numerical techniques. Based on the OGH, the recursion method has been applied in various lines fo research [27]. Particularly it has been applied successfully in the context of correlation functions amounting to studies of dynamical quantities such as equilibration times [28, 29] and diffusion constants [30, 31]. The success in these applications triggered the question to address the full time evolution of correlation functions in this framework. Recently, in Ref. [32], an approach has been suggested to obtain correlation functions via an expansion in terms of pseudomodes, or Ruelle-Pollicot resonances [33, 34]. This method however presupposes the exact form of the Lanczos coefficients beyond the numerical scope, in performing a linear continuation of the latter motivated by the OGH.
In our work we put forward an approach to compute correlation functions on the basis of only the numerically accessible Lanczos coefficients that is both remarkably accurate yet very cheap. We provide analytical arguments that suggest a generic applicability of our method in systems in accord with the OGH. We augment these arguments by numerical investigations of pertinent toy models and specific quantum many-body spin chains, and benchmark using state-of-the-art numerical methods.
@@
$\varphi_{n}(t):=i^{-n}(\mathcal{O}_{n}|\mathcal{O}(t)),$ (1)
@@
$\dot{\vec{\boldsymbol{\varphi}}}=\textbf{L}\cdot\vec{\boldsymbol{\varphi}},\quad\text{where}\ \textbf{L}_{nm}=b_{n}\delta_{n,m-1}-b_{n-1}\delta_{n,m+1}.$ (3)
@@
Imposing the initial condition $\varphi_{n}(0)=\delta_{n,0}$, solving Eq. (3) amounts to the computation of the infinite-temperature autocorrelation function of the observable of interest, as $\mathcal{C}(t)=(\mathcal{O}_{0}(t)|\mathcal{O}_{0})=\text{tr}\{\mathcal{O}(t)\mathcal{O}\}/\text{tr}\{\mathbb{1}\}=\varphi_{0}(t)$. Conversely, other entries of $\vec{\boldsymbol{\varphi}}$ represent correlation functions between the initial operator and other operators in Krylov space, cf. Eq. (1).
@@
In principle, solving the matrix equation (3) allows for an exact computation of dynamical quantities such as the autocorrelation function, given an infinite amount of Lanczos coefficients. However, typically only a low, two-digit number of $b_{n}$ is numerically feasible, depending on the structure of system and observable alike. To cope with this limitation, the Lanczos coefficients are sometimes extrapolated in accord with the operator growth hypothesis (OGH) [25]. The latter states that, for few-body observables in chaotic quantum many-body systems that do not have overlap with any conserved quantity, the Lanczos coefficients asymptotically grow linear. Numerically, this behavior has been affirmed in several relevant settings, see e.g. [25, 35, 36, 30, 31]. It is worth noticing, however, that any truncation of L, albeit at a high dimension, will give rise to echoes, i.e. unphysical revivals of the correlation function, see e.g. [30].
@@
$\mathcal{A}_{r}=\frac{1}{b_{1}}\prod_{n=1}^{r-1}\left(\frac{b_{n}}{b_{n+1}}\right)^{(-1)^{n}}\sqrt{\left(\frac{b_{r}}{b_{r+1}}\right)^{(-1)^{r}}}$ (4)
This estimate has been suggested in Ref. [37]. It is expected to yield better results with increasing $r$. Let $R$ denote the position after which Eq. (4) has converged suitably. Note that this involves $R+1$ Lanczos coefficients. Next, we infer $D$ as the smallest integer such that $\left(b_{1}/b_{R+1}\right)^{D}\leq\delta$, with some suitably chosen, small $\delta$ (below we employ $\delta=1/5$). This second condition requires the Lanczos coefficients to grow. With both $R,D$ and $R+D<n_{\max}$ we define
$s_{n}=\begin{cases}0&0\leq n\leq R-2\\ (-1)^{n+R+1}\binom{D}{n-R-1}&R-1\leq n\leq R+D-1.\end{cases}$ (5)
For reasons that shall be explained later, we refer to $D$ as the ”order” of the LP curve. If both above conditions, involving $R$ and $D$, respectively, are met within the accessible $b_{n}$, we expect the LP approach to yield rather accurate results. This conforms to all examples presented below. Note that the non-zero entries in Eq. (5) are the Pascal numbers with alternating sign, motivating the name of our scheme.
$\dot{\vec{\boldsymbol{\psi}}}=\textbf{S}\cdot\vec{\boldsymbol{\psi}},$ (6)
where $\vec{\boldsymbol{\psi}}(t)=(\psi_{0}(t),\psi_{1}(t),\dots,\psi_{N-1}(t))^{T}$ with $N:=R+D$ and
$\textbf{S}=\begin{pmatrix}0&-b_{1}&0&0&\dots\\ b_{1}&0&-b_{2}&0&\dots\\ 0&b_{2}&0&\ddots&\dots\\ \vdots&\vdots&\ddots&\ddots&-b_{N-1}\\ b_{N}s_{0}&\dots&b_{N}s_{N-3}&b_{N}s_{N-2}+b_{N-1}&b_{N}s_{N-1}\end{pmatrix},$ (7)
and the initial condition $\psi_{n}(0)=\delta_{n0}$. Similar approaches to truncations of the Liouvillian have been introduced in [38, 39]. The LP approach now asserts that $\psi_{n}(t)\approx\varphi_{n}(t)$ for $n=0,\dots,N-1$. Consequently, in analogy to Eq. (1), solutions of the simple $N\times N$ matrix equation (6) give rise to approximations of several relevant dynamical quantities ranging from autocorrelations to expectation values in quench settings. Concretely, we have
$\psi_{0}(t)$ LP${\approx}\left(\mathcal{O}_{0}(t)|\mathcal{O}_{0}\right)\propto\text{Tr}\{\mathcal{O}_{0}(t)\mathcal{O}_{0}\},$ (8)
$\psi_{1}(t)$ LP${\approx}\left(\mathcal{O}_{1}(t)|\mathcal{O}_{0}\right)\propto\text{Tr}\{\left[\mathcal{H},\mathcal{O}_{0}\right](t)\mathcal{O}_{0}\}\dots$ (9)
Numerics. Before turning to physical models, we consider settings in which the connection between the Lanczos coefficients and the analytical form of the autocorrelation function is known. To this end, we study the toy model given by $b_{n}=1+\alpha(n-1)$ for $0\leq\alpha\leq 1$. In the cases where $\alpha=0$ and $\alpha=1$, the autocorrelation functions are known to be $C(t)=J_{1}(2t)/t$ and $C(t)=\text{sech}(t)$ respectively [37]. Here $J_{n}(x)$ denotes the Bessel functions of first kind. The long-time behavior of the former is known to follow a power-law, i.e. an approximation with few exponentials is inherently flawed. We note that in line with our construction, for constant $b_{n}$, no $D$, as described below Eq. (4) can be inferred. Here we focus on the two cases $\alpha=0.1,1$ specifically. While for $\alpha=1$ we may compare with the analytical form of the autocorrelation function, for $\alpha=0.1$ we benchmark the LP scheme against the solution of the matrix equation (3) with dimension 2000  [^1] As explained in the introduction, solutions of Eq. (3) inevitably features echoes. Here we therefore only consider times where the correlation function has decayed suitably but well before the echo sets in..
Throughout the paper, we determine $R$ as the position after which $\mathcal{A}_{r}$ varies less that $1/200\max_{j}\mathcal{A}_{j}$ from subsequent areas. Further, we set $\delta=0.2$. In Fig. 1 we show the LP scheme at the example of the toy model. Here, the area $\mathcal{A}_{R}$ has converged in the above sense at $R=1$ for $\alpha=0.1$ and at $R=9$ for $\alpha=1$. The quantities $D$ are determined to be 19 and 1, respectively. We find that in both cases the LP method gives rise to approximations to the autocorrelation functions with strong agreement to the actual dynamics. For the toy model with $\alpha=0.1$ we also depict the respective LP curve with $(R,D)=(19,1)$ in order for a fair comparison such that in both scenarios the curves consist of the same number of damped oscillations. Unlike for $(R,D)=(1,19)$, the curve eventually fails to capture the true dynamics and deviates visibly indicating the importance of a suitable order in the LP scheme.
@@
Here, we consider up to $n_{\text{max}}=40$ Lanczos coefficients. See App. C for further examples in the tilted-field Ising model. Next, we include into our study a two-dimensional transverse-field Ising model,
@@
where by the notation $\langle ij\rangle$ the sum is restricted to adjacent sites on the two-dimensional lattice. As an observable of interest we consider the energy current
$\mathcal{J}_{E}^{2d}=\sum_{\langle ij\rangle|i<j}\sigma_{i}^{x}\sigma_{j}^{y}-\sigma_{j}^{x}\sigma_{i}^{y}.$ (13)
This setting has recently been studied in a related context, from where also the first 21 Lanczos coefficients are taken [32].
@@
In Fig. 2 we depict the autocorrelation functions for the models and observables introduced above. We find that whenever the LP approach is feasible, i.e. the area $\mathcal{A}_{R}$ has converged in the above sense and we may infer $D$, the LP curve is practically indistinguishable from the actual data. It is worth mentioning that this agreement is present, regardless of the feature of the autocorrelation functions, i.e. few or many oscillations or mere decay. Deviations set in on a scale that is negligible for most practical purposes. For the tilted-field Ising model with a strong perpendicular field $B_{z}$, here $B_{z}=2.0$, the areas $\mathcal{A}_{r}$ do not converge within our numerical scope, hence we cannot infer an LP curve subject to the construction introduced above and our $n_{\max}$. It is worth noticing that, for a magnetic field of this strength, the finite system leaves the chaotic regime, see e.g. [29]. To highlight the inapplicablity of the LP method in this scenario, we depict in Fig. 2 $m)$ two curves, each consisting of the maximal number of damped oscillations allowed by the LP scheme. Concretely, we show cases where $(R,D)=(40,1)$ and $(31,10)$ respectively.
Further, via the LP approach we may address other correlation functions. For the previous examples with convergent areas, we show in Fig. 3 the LP approximations $\psi_{1}(t)$ to higher correlators. As with the autocorrelator and $\psi_{0}(t)$ respectively, the agreement of the LP method is remarkable with deviations only visible on the log-scale.
For all observables that feature smoothly increasing Lanczos coefficients $b_{n}$, see Figs. 1 and 2 $[(c),(f),(i),(l)]$, we find the area estimator $\mathcal{A}_{r}$ to converge within our numerical scope and eventually yielding remarkably accurate approximations via the LP method. Conversely, for cases with ragged $b_{n}$ such as for the fast-mode observable in the TFIM with strong field, see Fig. 2 $(o)$, no LP curve may be obtained for our $n_{\max}$. Applicability of the Lanczos-Pascal method. The applicability of the LP scheme is closely tied to the structure of the Lanczos coefficients $b_{n}$ themselves, see also App. A,B. In order for the LP scheme to apply, the respective areas need to converge. Considering
@@
we infer that for convergent areas $\mathcal{A}_{r}$, the Lanczos coefficients need to become sufficiently smooth, i.e. $b_{r}b_{r+2}/b_{r+1}^{2}\approx 1$. It is worth stressing that for systems within the scope of the operator growth hypothesis, the latter condition may be expected to be generically fulfilled. Revisiting the Lanczos coefficients for the settings considered in this paper in Fig. 2 and Fig. 3, precisely the cases with smoothly increasing $b_{n}$ each give rise to accurate descriptions by means of the LP method. In essence, the comparatively weak requirement posed in Eq. (15) strongly suggests the generic applicability of the LP method in chaotic quantum many-body systems.
Conclusion. In this paper, we introduced the Lanczos-Pascal approach as a scheme based on the recursion method to approximate dynamics in quantum many-body systems. Taking into account only few, numerically accessible Lanczos coefficients, the method is very cheap. We formulated criteria based on the Lanczos coefficients of an observable in order to a priori assess whether the proposed method is applicable in a given setting. We numerically checked the method for a multitude of systems, ranging from a toy model, over pertinent chaotic Ising chains in both one and two dimension to an open system spin-bath setup. In all cases, the LP approximations agree remarkably well with state-of-the-art benchmarks obtained by means of dynamical quantum typicality, given that predefined conditions on the Lanczos coefficients indicated the LP to be feasible.
We augment our numerical findings by analytical considerations that link the convergence of the LP scheme to structures in the respective Lanczos coefficients. Crucially, in the context of few-body observables in chaotic quantum many-body systems, these considerations are supplemented by the operator growth hypothesis (OGH). With the latter putting forward an eventual linear growth of the Lanczos coefficients, our arguments suggest the generic applicability of our scheme. In App. B we provide similar arguments for the case of finite frequencies $\omega\neq 0.$
Moreover, the feasibility of the LP approach has further conceptual implications. By construction the LP approach involving $m$ Lanczos coefficients will yield approximations that each consist of at most $m$ damped exponentials. In this interpretation, curves that yield LP curves with low $m$ may be regarded as simple. This addresses nature’s apparent soft spot for simple dynamics, oftentimes appearing to favor certain classes of relaxation over others.
@@
Appendix A The Lanczos-Pascal scheme as a derivative in Krylov space

### v3

@@ Abstract
Introduction. Understanding out-of-equilibrium quantum dynamics is one of the most fundamental and challenging questions in the field of statistical physics. It has received significant attention, from both theoretical and experimental perspectives [1, 2, 3, 4]. Despite the substantial progress that has been made, many fundamental aspects of the problem remain unresolved, and it continues to be the focus of intensive research in the field.
@@
To characterize the evolution of correlation functions, on the analytical side, general predictions remain a major difficulty, except for specific cases such as free fermions and certain nontrivial but analytically accessible cases, including particular quantum circuits [9, 10, 11]. In systems with conserved charge, hydrodynamics has been proven to be a useful tool to predict the long-time behavior of autocorrelation functions [12, 13, 14]. However, it remains an assumption, and the extent of its applicability has yet to be fully established.
@@
Another approach, the recursion method [25, 26, 27], significantly regained attention thanks to the recent Operator Growth Hypothesis (OGH) [28]. The latter constitutes a statement on the so-called Lanczos coefficients of an observable; quantities that fully characterize its dynamics in the thermodynamic limit. Eventually code bases such as [29] position this method along other well-established numerical techniques. Based on the OGH, the recursion method has been applied in various lines fo research [30]. Particularly it has been applied successfully in the context of correlation functions amounting to studies of dynamical quantities such as equilibration times [31, 32] and diffusion constants [33, 34]. The success in these applications triggered the question to address the full time evolution of correlation functions in this framework. Recently, in Ref. [35], an approach has been suggested to obtain correlation functions via an expansion in terms of pseudomodes, or Ruelle-Pollicott resonances [36, 37], by relying on an extrapolation of the Lanczos coefficients and considering a vanishing artificial dissipation. Besides these endeavors based on the Lanczos coefficients, progress has been made in the computation of correlation functions of hydrodynamical observables including procedures similar to truncated Lanczos expansions, based on tensor networks [38, 39, 40].
In our work we put forward an approach to directly compute correlation functions on the basis of only the practically numerically accessible Lanczos coefficients, i.e. without further explicit assumption on numerical inaccessible ones. The approach is both remarkably accurate yet very cheap  [^1] While the cost to compute Lanczos coefficients scales exponentially in the number of coefficients, the first $\sim 20-30$ Lanczos coefficients (depending on the model) are readily computable. As will become evident throughout the paper, this number of coefficients typically suffices to apply the scheme.. We provide analytical arguments that establish the applicability of our method in settings with smooth Lanczos coefficients in accord with the OGH. We augment these arguments by numerical investigations of pertinent toy models and specific quantum many-body spin chains, and benchmark using state-of-the-art numerical methods.
@@
$\varphi_{n}(t):=i^{-n}(\mathcal{O}_{n}|\mathcal{O}_{0}(t)),$ (1)
@@
$\dot{\vec{\boldsymbol{\varphi}}}=\textbf{L}\cdot\vec{\boldsymbol{\varphi}},\quad\text{where}\ \textbf{L}_{nm}=b_{n-1}\delta_{n,m+1}-b_{n}\delta_{n,m-1}.$ (3)
@@
Imposing the initial condition $\varphi_{n}(0)=\delta_{n,0}$, solving Eq. (3) amounts to the computation of the infinite-temperature autocorrelation function of the observable of interest, as $\mathcal{C}(t)=(\mathcal{O}_{0}(t)|\mathcal{O}_{0})=\text{tr}\{\mathcal{O}_{0}(t)\mathcal{O}_{0}\}/\text{tr}\{\mathbb{1}\}=\varphi_{0}(t)$. Conversely, other entries of $\vec{\boldsymbol{\varphi}}$ represent correlation functions between the initial operator and other operators in Krylov space, cf. Eq. (1). Note that, for condensed-matter-type settings, frequently $\mathcal{A}$ and $\mathcal{H}$ are local. In this case the first $fL$ coefficients $b_{n}$ do not depend on system size $L$, with $f$ being a number of the order of one [27].
@@
In principle, solving the matrix equation (3) allows for an exact computation of dynamical quantities such as the autocorrelation function, given an infinite amount of Lanczos coefficients. However, typically only a low, two-digit number of $b_{n}$ pertaining to the infinite system is numerically feasible, depending on the structure of system and observable alike. To cope with this limitation, the Lanczos coefficients are sometimes extrapolated in accord with the operator growth hypothesis (OGH) [28]. The latter states that, for few-body observables in chaotic quantum many-body systems that do not have overlap with any conserved quantity, the Lanczos coefficients asymptotically grow linear. Numerically, this behavior has been affirmed in several relevant settings, see e.g. [28, 42, 43, 33, 34]. It is worth noticing, however, that any truncation of L, albeit at a high dimension, will give rise to echoes, i.e. unphysical revivals of the correlation function, see e.g. [33].
Lanczos-Pascal approach. Concretely, the Lanczos-Pascal (LP) approach (as a recipe) proceeds as follows. ( Derivations of the applicability of the LP approach under the below-mentioned conditions on the $b_{n}$ may be found in End Matter (EM) A,B.) As a first step, we determine an area estimation $\mathcal{A}_{r}$ for $\varphi_{0}(t)$, i.e., $\mathcal{A}_{r}\approx\int_{0}^{\infty}\varphi_{0}(t)$dtt on the basis of the available Lanczos coefficients, given by
$\mathcal{A}_{r}=\frac{1}{b_{1}}\prod_{n=1}^{r-1}\left(\frac{b_{n}}{b_{n+1}}\right)^{(-1)^{n}}\sqrt{\left(\frac{b_{r}}{b_{r+1}}\right)^{(-1)^{r}}}.$ (4)
This estimate has been suggested in Ref. [44]. It is expected to yield better results with increasing $r$. Let $R$ denote the position after which Eq. (4) has converged suitably. Note that this involves $R+1$ Lanczos coefficients. Next, we infer $D$ as the smallest integer such that $\left(b_{1}/b_{R+1}\right)^{D}\leq\delta$, with some suitably chosen, small $\delta$ (below we employ $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\delta=1/10$}). This second condition requires the Lanczos coefficients to grow. With both $R,D$ and $R+D<n_{\max}$ we define
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s_{n}}$ $=\begin{cases}0&0\leq n\leq R-1\\ {\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(-1)^{n+R+D}\binom{D}{n-R}}&{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}R\leq n\leq R+D-1}.\end{cases}$ (5)
For reasons that shall be explained later, we refer to $D$ as the ”order” of the LP curve. If both above conditions, involving $R$ and $D$, respectively, are met within the accessible $b_{n}$, we expect the LP approach to yield rather accurate results. This conforms to all examples presented below. Note that the nonzero entries in Eq. (5) are the Pascal numbers with alternating sign, motivating the name of our scheme.
$\dot{\vec{\boldsymbol{\psi}}}=\textbf{S}\cdot\vec{\boldsymbol{\psi}},\quad\text{where}\ {\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{S}_{nm}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}=\textbf{L}_{nm}+\delta_{n,N-1}b_{N}s_{m}}$ (6)
with $\vec{\boldsymbol{\psi}}(t)=(\psi_{0}(t),\psi_{1}(t),\dots,\psi_{N-1}(t))^{T}$ and $N:=R+D$ and the initial condition $\psi_{n}(0)=\delta_{n0}$. Similar approaches to truncations of the Liouvillian have been introduced in [45, 46]. The LP approach now asserts that $\psi_{n}(t)\approx\varphi_{n}(t)$ for $n=0,\dots,N-1$. Consequently, in analogy to Eq. (1), solutions of the simple $N\times N$ matrix equation (6) give rise to approximations of several relevant dynamical quantities ranging from autocorrelations to expectation values in quench settings. Concretely, we have
$\psi_{0}(t)$ $\underset{\text{LP}}{\approx}\left(\mathcal{O}_{0}(t)|\mathcal{O}_{0}\right)\propto\text{Tr}\{\mathcal{O}_{0}(t)\mathcal{O}_{0}\},$ (7)
$\psi_{1}(t)$ $\underset{\text{LP}}{\approx}\left(\mathcal{O}_{1}(t)|\mathcal{O}_{0}\right)\propto\text{Tr}\{\left[\mathcal{H},\mathcal{O}_{0}\right](t)\mathcal{O}_{0}\}\dots$ (8)
@@
Throughout the paper, we determine $R$ as the position after which $\mathcal{A}_{r}$ varies less that $1/200\max_{j}\mathcal{A}_{j}$ from subsequent areas. In Fig. 1 we show the LP scheme at the example of the toy model. Here, the area $\mathcal{A}_{R}$ has converged in the above sense at $R=1$ for $\alpha=0.1$ and at $R=9$ for $\alpha=1$. The quantities $D$ are determined to be 27 and 1, respectively. We find that in both cases the LP method gives rise to approximations to the autocorrelation functions with strong agreement to the actual dynamics. For the toy model with $\alpha=0.1$ we also depict the respective LP curve with $(R,D)=({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}27},1)$ in order for a fair comparison such that in both scenarios the curves consist of the same number of damped oscillations. Unlike for $(R,D)=(1,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}27})$, the curve eventually fails to capture the true dynamics and deviates visibly indicating the importance of a suitable order in the LP scheme.
@@
Here, we consider up to $n_{\text{max}}=40$ Lanczos coefficients. See the Supplemental Material [48] for further examples in the tilted-field Ising model. Next, we include into our study a two-dimensional transverse-field Ising model,
@@
where by the notation $\langle ij\rangle$ the sum is restricted to adjacent sites on the two-dimensional lattice. As an observable of interest we consider the horizontal energy current
$\mathcal{J}_{E}^{2d}=\sum_{\langle ij\rangle|i<j}\sigma_{i}^{x}\sigma_{j}^{y}-\sigma_{j}^{x}\sigma_{i}^{y},$ (12)
with $i<j$ restricting the sum to horizontally neighboring sites. This setting has recently been studied in a related context, from where also the first 21 Lanczos coefficients are taken [35].
@@
In Fig. 2 we depict the autocorrelation functions for the models and observables introduced above. We find that whenever the LP approach is feasible, i.e. the area $\mathcal{A}_{R}$ has converged in the above sense and we may infer $D$, the LP curve is practically indistinguishable from the actual data. It is worth mentioning that this agreement is present, regardless of the feature of the autocorrelation functions, i.e. few or many oscillations or mere decay. Deviations set in on a scale that is negligible for most practical purposes. For the tilted-field Ising model with a strong perpendicular field $B_{z}$, here $B_{z}=2.0$, the areas $\mathcal{A}_{r}$ do not converge within our numerical scope, hence we cannot infer an LP curve subject to the construction introduced above and our $n_{\max}$. Note that, for a magnetic field of this strength, the finite system leaves the chaotic regime, see e.g. [32]. To highlight the inapplicablity of the LP method in this scenario, we depict in Fig. 2 $m)$ two curves, each consisting of the maximal number of damped oscillations allowed by the number of feasible Lanczos coefficients. Concretely, we show cases where $(R,D)=(40,1)$ and $(31,10)$ respectively [^3] Note that the parameters $R,D$ for this setting only serve to illustrate the dynamics. They are not inferred with respect to the convergence criteria as described in the main text..
Further, via the LP approach we may address other correlation functions.
For the previous examples with convergent areas, we show in Fig. 3 the LP approximations $\psi_{1}(t)$ to higher correlators. As with the autocorrelator and $\psi_{0}(t)$ respectively, the agreement of the LP method is remarkable with deviations only visible on the log-scale.
For all observables that feature smoothly increasing Lanczos coefficients $b_{n}$, see Figs. 1 and 2 $[(c),(f),(i),(l)]$, we find the area estimator $\mathcal{A}_{r}$ to converge within our numerical scope and eventually yielding remarkably accurate approximations via the LP method. Conversely, for cases with ragged $b_{n}$ such as for the fast-mode observable in the TFIM with strong field, see Fig. 2 $(o)$, no LP curve may be obtained for our $n_{\max}$.
Applicability of the Lanczos-Pascal method. The applicability of the LP scheme is closely tied to the structure of the Lanczos coefficients $b_{n}$ themselves, see also EM B. In order for the LP scheme to apply, the respective areas need to converge. Considering
@@
we infer that for convergent areas $\mathcal{A}_{r}$, the Lanczos coefficients need to become sufficiently smooth, i.e. $b_{r}b_{r+2}/b_{r+1}^{2}\approx 1$  [^4] A similar analysis in a related context has already been discussed in Ref. [44]. Revisiting the Lanczos coefficients for the settings considered in this paper in Fig. 2 and Fig. 3, precisely the cases with smoothly increasing $b_{n}$ each give rise to accurate descriptions by means of the LP method. Moreover, the smoothness of the increase appears to be present in the vast majority of scenarios, while features causing (exceedingly) high $R$ or $D$ such as staggering or absence or delay of growth of the $b_{n}$, are mainly found in settings outside the scope of the OGH, i.e. for observables overlapping with conserved quantities or systems away from full chaoticity, see [28, 32, 34, 33, 35, 42, 52, 43].
Conclusion and Outlook. We introduced the Lanczos-Pascal approach as a scheme based on the recursion method to approximate high temperature autocorrelation functions in quantum many-body systems in the thermodynamic limit. We present readily checkable criteria based on the respective Lanczos coefficients for the applicability of the LP scheme and find the latter fulfilled for a variety of examples featuring chaotic quantum systems and observables that have no overlap with a conserved quantity. We checked the results of the LP approach against data obtained from dynamical quantum typicality and found good agreement. DQT itself is known to produce results that are in good agreement with state-of-the-art techniques, e.g., based on matrix product states etc. [17]. We do not necessarily expect the LP approach to be more accurate or have a wider range of applicability than the above state-of-the-art techniques. The impact of the LP-approach rather roots in two features: $i)$ within the regime of its applicability the computational costs for the LP method are much lower than for other methods, ii)ii) without even performing a concrete calculation, the LP approach establishes that, again within the regime of its applicability, the respective correlation functions are well described by a low number of exponentially damped oscillations. The latter dynamics are ubiquitous in nature, yet to this date encompassing explanations for this ubiquity are scarce.
Conversely, in cases where the criteria are not met and hence the LP method becomes inapplicable, this suggests that the underlying correlation function exhibits more intricate behavior, e.g. as seen in [53, 54]. Exploring the possible extension of the LP approach to such cases is an interesting topic for future study.
@@
Data availability. The data that support the findings of this article are openly available [55].
Appendix A Applicability of the Lanczos-Pascal scheme under the assumption of smooth $\varphi_{n}(t)$